# Belief Bias Identification

### v2

@@ 6 Concluding remarks
In conclusion, this paper addresses the challenges in identifying and testing multiple belief-updating biases simultaneously. This is important because in the absence of such a model, many of these biases could be potentially confounded. The theoretical model introduces a novel approach by considering belief distributions over a continuous state space, allowing for a complete identification of multiple conflicting belief biases. The corresponding laboratory experiment applies this model, using a novel technique to elicit belief distributions.
The population-level analysis shows that, despite the prevalence of overinference and base rate neglect in the baseline model, only base rate neglect remains significant in the complete model, emphasizing the importance of accounting for a broader spectrum of biases. Individual-level analysis uncovers significant heterogeneity, with every participant exhibiting at least one identifiable bias. Further exploration into biases’ impact on expected value and variance exposes motivated-belief biases (optimism and pessimism) and sequence-related biases (gambler’s and hot hand fallacies) as key drivers of biased inference. Thus, capturing a wide array of biases provides a much more subtle understanding of belief updating dynamics. Overall, addressing confoundedness in belief-updating biases is key to researchers interested in understanding links between different biases, or delving deeper into the causes of various behavioral phenomena (political polarization, investing performance, disposition effect…) when conflicting updating biases are at stake.
I would like to conclude by pointing to some related future research avenues. While this paper takes detailed care in separating behavior stemming from different biases, it is important to note that it approaches biases as systematic departures from rational behavior. That is, it does not dig into the cognitive mechanisms or channels by which a certain bias might be exhibited. In this direction Bordalo et al., 2023a , look at how some biases become more prominent when salient features of the information become more relevant. At the same time, Bordalo et al., 2023b also emphasize the importance of memory and recall to explain well-known cognitive biases. Further exploring more of these cognitive mechanisms in a setting where biases can be de-confounded, could deliver a complete picture of what updating biases matter the most, and pinpoint the primitive traits that underlie biased behavior.
Appendix to Belief Bias Identification
Appendix A Conjugate analyses of the binomial distribution
@@
This section shows the derivation of equations of posterior parameters in section 2. I present two cases: The Bayesian case (eq. (3)), and introducing the biases of the baseline model (eq.(4) and (5)). The rest of the proofs for the remaining biases are analogous to these cases. Once the distorted likelihood and prior are given, the proof follows the exact same steps.
A.1 The Bayesian case
Statement 1.
@@ A.1 The Bayesian case
Which is itself the probability density function of a beta distribution with parameters ($a_{0}+k,b_{0}+n-k$). Therefore, the posterior Bayesian distribution is beta distributed with parameters ($a_{n},b_{n}$) as defined in Statement 1.
A.2 Biases of the baseline model: Inference and base-rate biases
Statement 2.
@@ A.2 Biases of the baseline model: Inference and base-rate biases
[…]
@@
This is the probability density function of a beta distribution with parameters $\Big{(}\gamma k+\delta(a_{0}-1)+1,\gamma(n-k)+\delta(b_{0}-1)+1\Big{)}$. Therefore, the posterior distribution of an agent that exhibits inference bias $\gamma$, and base-rate bias ($\delta$) is beta distributed with parameters ($\tilde{a}_{n},\tilde{b}_{n}$) as defined in Statement 2. ∎
Appendix B Incentivising mean and variance of beta distributions
The scoring rules follow Schlag et al., (2015) with a slight modification. In particular, random realisations of the Bayesian posterior distributions are taken as the random draws.
B.1 Incentivising the mean
@@ C.1 Information criteria and $R^{2}$ comparisons
Table 4: $R^{2}$ Comparison for Baseline and Complete model at population vs individual level
C.2 Individual bias frequency corrected for multiple hypothesis testing
As the individual-level analysis in section 5.2 tests a considerable number of null hypotheses (eleven per subject), one may wonder whether some of these results are driven by a high rate of false positives. To address this concern, this section replicates Figures 4(a) and 4(b) after applying the Šidák correction for multiple hypothesis testing. In particular, $p<1-(1-0.05)^{1/11}\approx 0.0047$
@@ C.2 Individual bias frequency corrected for multiple hypothesis testing
(a) Figure 7a: Number of times a specific bias is found to be significant ($p<0.047$) in the baseline model at the individual level.
@@
(b) Figure 7b: Number of times a specific bias is found to be significant ($p<0.047$) in the complete model at the individual level.
As one can see from figures 7(a) and 7(b), the three main results from section 5.2 still hold. Namely, i) all biases are exhibited to a certain extent in the complete model, ii) transitioning from the baseline to the complete model shows that all subjects exhibit some kind of bias (empty “No Bias” column in figure 7(b)), and iii) the hot hand effect stands out as the most-frequently exhibited bias.
In fact, the second result seems to be even stronger under this condition, with the vast majority of subjects appearing to be “too noisy” in the baseline model, but showing some type of bias in the complete model. However, the relative frequency of motivated-belief biases (optimism/pessimism) significantly diminishes under $p<0.0047$.
Appendix D Experiment instructions

### v4

@@ 6 Concluding remarks
This paper develops a unified framework for separately identifying multiple belief-updating biases that may otherwise be empirically confounded. The key methodological idea is to study belief updating using prior and posterior belief distributions over a continuous state space, rather than point beliefs alone. This allows a broader range of conflicting distortions to be brought into the same structural framework and tested in a laboratory setting using a novel belief-elicitation method.
The results show that accounting for a richer set of biases changes the interpretation of updating behavior in important ways. At the population level, overinference and base-rate neglect are both significant in the baseline specification, but once the complete model is estimated, only base-rate neglect remains significant. At the individual level, the analysis uncovers substantial heterogeneity: under the complete specification, all tested bias types are present to some extent in the data, and more subjects are categorized as exhibiting some type of bias. Motivated-belief biases and sequence-related biases emerge as the main drivers of biased inference, while confirmatory distortions are comparatively scarce. Moreover, biases do not arise independently. Rather, they exhibit systematic co-occurrence patterns that suggest a small number of recurring configurations, with sequence-related distortions appearing central in the overall structure.
Taken together, these findings suggest that studying belief-updating biases one at a time may give an incomplete picture of the underlying mechanisms shaping inference. More generally, a framework that separates multiple competing distortions can be useful for researchers interested in the interaction between biases, as well as for work seeking belief-based explanations of broader behavioral phenomena, such as political polarization, investment behavior, or the disposition effect.
Several directions for future research remain open. While this paper takes detailed care in separating behavior stemming from different biases, it treats such biases as systematic departures from Bayesian updating and does not attempt to identify the deeper cognitive mechanisms through which they arise. In this direction, Bordalo et al. (2025) study how some biases become more prominent when salient features of information become more relevant. Likewise, Bordalo et al. (2023) emphasize the role of memory and recall in generating well-known cognitive distortions, while Graeber et al. (2024) show how different forms of information –such as narratives versus statistical evidence– affect belief updating. Exploring these mechanisms in settings where biases can already be separately identified may help provide a more complete account of which updating distortions matter most, and of the underlying primitives that generate biased behavior.
Appendix
Appendix A Theoretical derivations
@@
This section shows the derivation of equations of posterior parameters in section 2. I present two cases: Proposition 1 characterizes the Bayesian benchmark under standard beta-binomial conjugacy (eq.(3)), and Proposition 2 introduces the biases of the baseline model (eq.(4) and (5)). The rest of the derivations for the remaining biases are analogous to these cases. Once the distorted likelihood and prior are given, the proof follows the exact same steps.
Proposition 1.
@@
Which is itself the probability density function of a beta distribution with parameters ($a_{0}+k,b_{0}+n-k$). Therefore, the posterior Bayesian distribution is beta distributed with parameters ($a_{n},b_{n}$) as defined in Proposition 1.
Proposition 2.
@@
$\tilde{\pi}(p|s_{1}...s_{n},a_{0},b_{0})$ $=\frac{(L(p|s_{1}...s_{n}))^{\gamma}(\pi(p))^{\delta}}{\int_{p=0}^{1}\big((L(p|s_{1}...s_{n}))^{\gamma}(\pi(p))^{\delta}\big)dp}$
$=\frac{\big[\binom{n}{k}\big]^{\gamma}p^{\gamma k}(1-p)^{\gamma(n-k)}\big(1/B(a_{0},b_{0})\big)^{\delta}p^{\delta(a_{0}-1}(1-p)^{\delta(b_{0}-1)}}{\int_{p=0}^{1}\Big(\big[\binom{n}{k}\big]^{\gamma}p^{\gamma k}(1-p)^{\gamma(n-k)}\big(1/B(a_{0},b_{0})\big)^{\delta}p^{\delta(a_{0}-1}(1-p)^{\delta(b_{0}-1)}\Big)dp}$
$=\frac{p^{\gamma k+\delta(a_{0}-1)}(1-p)^{\gamma(n-k)+\delta(b_{0}-1)}}{\int_{p=0}^{1}\Big(p^{\gamma k+\delta(a_{0}-1)}(1-p)^{\gamma(n-k)+\delta(b_{0}-1)}\Big)dp}$
@@
This is the probability density function of a beta distribution with parameters $\Big(\gamma k+\delta(a_{0}-1)+1,\gamma(n-k)+\delta(b_{0}-1)+1\Big)$. Therefore, the posterior distribution of an agent that exhibits inference bias $\gamma$, and base-rate bias ($\delta$) is beta distributed with parameters ($\tilde{a}_{n},\tilde{b}_{n}$) as defined in Proposition 2. ∎
Appendix B Incentivizing moments of beta distributions
The scoring rules follow Schlag and Van der Weele (2013) with a slight modification. In particular, random realizations of the Bayesian posterior distributions are taken as the random draws.
Incentivizing the mean.
@@ Appendix C Extra tables and figures
Table 4: $R^{2}$ Comparison for Baseline and Complete model at population vs. individual level
Appendix D Robustness analyses
D.1 Robustness to Measurement Error in Uncertainty Reports
Uncertainty is elicited through a continuous slider and may therefore be measured with classical measurement error. To assess whether such measurement error affects the bias classification, I borrow the noise-injection logic from the SIMEX (Simulation Extrapolation) method Cook and Stefanski (1994) for parametric models, and apply it to both the prior and posterior reported variance.
Let $\mu$ denote the reported belief mean (be it prior or posterior) and $V$ the reported variance. Because beliefs are defined on the bounded support $[0,1]$, reported variances must satisfy
$0<V<V_{\max}(\mu),\qquad V_{\max}(\mu)=\mu(1-\mu)$
Note that this restriction follows from the support of the distribution and does not impose unimodality of the implied beta distribution; bimodal beta distributions remain feasible whenever the reported variance is sufficiently close to the upper bound. To model reporting noise while preserving feasibility, I work with the normalized variance ratio
$u=\frac{V}{V_{\max}(\mu)}\in(0,1)$
and apply the logit transformation $z=\log\!\left(\frac{u}{1-u}\right)\in\mathbb{R}$.
Measurement error is assumed to be classical and additive on this scale:
$z^{\text{rep}}=z^{\star}+\sigma_{z}\varepsilon\qquad\varepsilon\sim\mathcal{N}(0,1)$
where $z^{\star}$ denotes the latent (noise-free) report and $\sigma_{z}$ is the standard deviation of reporting noise on the logit-normalized variance scale.
To implement this robustness exercise, I simulate additional noise on the logit-normalized variance scale. For each $\lambda\in\Lambda=\{0.1,0.25,0.5\}$ and each observation, the perturbed report is given by
$z^{(\lambda)}=z^{\text{rep}}+\sqrt{\lambda}\,\sigma_{z}\,\varepsilon^{(\lambda)}\qquad\varepsilon^{(\lambda)}\sim\mathcal{N}(0,1)$ (24)
where $\lambda$ indexes the severity of the perturbation. Thus, $\lambda=0.1$, $\lambda=0.25$, and $\lambda=0.5$ correspond to adding noise with variance equal to $10\%$, $25\%$, and $50\%$ of the estimated baseline measurement-error variance, respectively. This range is intended to capture progressively stronger perturbations of the reported variance while remaining within a plausible range for slider misreporting.
A baseline value $\sigma_{z}=\hat{\sigma}_{z}$ is calibrated from the data as the residual standard deviation from a regression of reported logit-normalized variances on their Bayesian benchmark variances on the same scale. Mapping back via $u^{(\lambda)}=\text{logit}^{-1}(z^{(\lambda)})$ yields simulated feasible variances (both for prior and posterior measurements)
$V^{(\lambda)}=u^{(\lambda)}V_{\max}(\mu)$
For each $\lambda$, the simulation is repeated $200$ times and the adjusted variances $V^{(\lambda)}$ are averaged across repetitions. The procedure is applied to both prior and posterior variance reports. For each simulated variance, the implied beta-distribution parameters are recomputed and all model equations are re-estimated using these noise-adjusted values.
This section replicates Figures 4(a), 4(b), 5(b) and 6(b) in the main text for the largest simulated noise level, $\lambda=0.5$.
@@ D.1 Robustness to Measurement Error in Uncertainty Reports
(a) Figure 9a: Number of times a specific bias is found to be significant ($p<0.05$) in the baseline model at the individual level after noise injection of $\lambda=0.5$ in reported variances.
@@
(b) Figure 9b: Number of times a specific bias is found to be significant ($p<0.05$) in the complete model at the individual level after noise injection of $\lambda=0.5$ in reported variances.
As Figures 9(a) and 9(b) show, many of the main qualitative conclusions remain unchanged even under this conservative level of noise injection. Every bias continues to be present in the complete model, the complete model still categorizes a substantially larger share of subjects than the baseline model, and hot-hand effect remains the most commonly exhibited bias. Moreover, Figures 10(a) and 10(b) show that sequence-related biases (gambler’s fallacy and hot-hand behavior) and motivated beliefs (especially optimism) continue to account for most of the inference in expected-value and variance deviations. The main change is that overinference becomes less prevalent, while underinference becomes substantially more prominent.
(a) Figure 10a: Expected-value deviations with respect to the Bayesian framework (adjusted by bias prevalence) after noise injection of $\lambda=0.5$.
(b) Figure 10b: Variance deviations with respect to the Bayesian framework (adjusted by bias prevalence) after noise injection of $\lambda=0.5$.
D.2 Individual bias frequency corrected for multiple hypothesis testing
As the individual-level analysis in Section 5.2 classifies subjects on the basis of multiple coefficient tests, a natural concern is that some detected biases may reflect false positives. This concern is particularly relevant when comparing the baseline and complete models, since the baseline model tests four hypotheses per subject, whereas the complete model tests twelve. To address this issue, this appendix repeats the individual-level analysis after applying a Bonferroni correction within each model. Accordingly, the significance threshold is set to $p<0.05/4=0.0125$ for the baseline model, and to $p<0.05/12\approx 0.0042$ for the complete model. This section replicates Figures 4(a), 4(b), 5(b) and 6(b) under these corrected thresholds.
(a) Figure 11a: Number of times a specific bias is found to be significant ($p<0.0125$) in the baseline model at the individual level after multiple hypothesis testing correction.
(b) Figure 11b: Number of times a specific bias is found to be significant ($p<0.0042$) in the complete model at the individual level after multiple hypothesis testing correction.
(a) Figure 12a: Expected-value deviations with respect to the Bayesian framework (adjusted by bias prevalence at $p<0.0042$).
(b) Figure 12b: Variance deviations with respect to the Bayesian framework (adjusted by bias prevalence at $p<0.0042$)
The main qualitative conclusions remain largely unchanged: Every bias continues to be present in the complete model, and the hot-hand fallacy continues to be the most commonly exhibited bias (See Figures 11(a) and 11(b)). Also, sequence-related biases remain the main drivers of distorted inference, with gambler’s fallacy continuing to play a leading role for expected-value deviations and hot-hand behavior for variance deviations (See Figures 12(a) and 12(b)).
At the same time, the correction naturally reduces the number of detected significant biases and increases the number of unclassified subjects in both models. Nevertheless, the complete model still classifies substantially more subjects than the baseline model. In particular, the number of subjects in the “No Bias” category falls from $31$ under the baseline model to $20$ under the complete model (a $\approx 35\%$ reduction). Finally, it is important to note that, motivated-belief biases –especially optimism and pessimism– become relatively less prevalent under the corrected thresholds.
Appendix E A Grether-style reduced-form model
This section connects the main continuous state-space analysis to the classic reduced-form literature on belief updating in the spirit of Grether (1980). To do so, I approximate the continuous state space of the model by a binary state space and estimate a Grether-style regression using the experimental data. This exercise is intended to show that applying the usual models to this experimental data recovers patterns that can be related both to the classic literature and to the baseline findings of the main model.
Let the state space $\Theta=\{L,H\}$ be a binary partition of the original state-space $\Omega=(0,1)$, where state $H$ corresponds to higher success probabilities and state $L$ to lower success probabilities.
Let $q_{0}=\Pr(H)$ denote the prior belief assigned to the high state, and let $q_{n}=\Pr(H\mid k,n)$ denote the corresponding posterior belief after observing $k$ successes in $n$ Bernoulli trials. Conditional on state $\theta\in\{L,H\}$, the probability of observing $k$ successes is
$\Pr(k\mid\theta)=\int_{p\in\theta}\binom{n}{k}p^{k}(1-p)^{n-k}f(p\mid\theta)\,dp$
where $f(p\mid\theta)$ denotes the prior density over $p$ conditional on state $\theta$.
The associated log-likelihood ratio is therefore
$\ell(k,n)=\log\left(\frac{\Pr(k\mid H)}{\Pr(k\mid L)}\right)$
Under Bayesian updating, posterior log-odds satisfy
$\log\left(\frac{q_{n}}{1-q_{n}}\right)=\log\left(\frac{q_{0}}{1-q_{0}}\right)+\ell(k,n)$
A Grether-style reduced form allows the prior and the likelihood ratio to receive separate weights in the updating equation. I therefore estimate
$\log\left(\frac{\tilde{q}_{n}}{1-\tilde{q}_{n}}\right)=\beta_{0}+\beta_{P}\log\left(\frac{q_{0}}{1-q_{0}}\right)+\beta_{L}\,\ell(k,n)+\varepsilon$ (25)
where $\tilde{q}_{n}$ denotes the reported posterior probability of the high state. The Bayesian benchmark is given by
$\beta_{0}=0\qquad\beta_{P}=1\qquad\beta_{L}=1$
Values $\beta_{P}<1$ indicate base-rate neglect, while $\beta_{P}>1$ indicate base-rate overuse. Similarly, $\beta_{L}<1$ indicates underinference, $\beta_{L}>1$ indicates overinference, and $\beta_{L}<0$ corresponds to inference against the signal, as in classic Grether (1980).
To implement this reduced-form exercise, I binarize the continuous state space at $p=0.5$, so that the high state is $H=\{p\geq 0.5\}$ and the low state is $L=\{p<0.5\}$. For each elicited prior and posterior beta distribution, I compute the implied probability assigned to the high state, $q=\Pr(p\geq 0.5),$ and use these probabilities as the belief objects entering the Grether-style regression. The evidence term is computed as the Bayesian log-odds update implied by the observed signal under the same binary partition, $\log\!\left(\frac{q_{n}^{B}}{1-q_{n}^{B}}\right)-\log\!\left(\frac{q_{0}}{1-q_{0}}\right),$ which is equivalent to $\ell(k,n)$ under this partition, where $q_{0}$ is the prior probability of the high state and $q_{n}^{B}$ is the Bayesian posterior probability of that state.
Dependent variable:
Log-odds reported posterior
Prior log-odds 0.681∗∗∗
Bayesian log-likelihood ratio 2.196∗∗∗
Constant 0.303∗∗∗
Observations 2,635
R2 0.559
Adjusted R2 0.559
Note: p<<0.1; ∗∗p<<0.05; ∗∗∗p<<0.01
Table 5: Grether-style model at the population level. Significance is with respect to Bayesian values. Clustered standard errors by participant.
The Grether-style reduced-form estimates in Table 5 are first and foremost consistent with the baseline analysis of the paper. At the population level, the prior-weight coefficient lies below the Bayesian benchmark of one ($\hat{\beta}_{P}=0.681$), indicating base-rate neglect, while the evidence-weight coefficient lies above one ($\hat{\beta}_{L}=2.196$), indicating overinference on average. Thus, once the data are forced into a traditional binary-state reduced form, the main qualitative pattern resembles that of the baseline model: subjects appear to underweight prior information and overweight new evidence.
At the same time, these estimates remain meaningfully connected to the classic reduced-form literature on belief updating. Base-rate neglect is the standard qualitative finding in the Grether tradition, and the presence of overinference, while less canonical than underinference (Benjamin, 2019), is by no means implausible. More recent contributions also document environments in which subjects overweight new signals rather than underreact to them (Augenblick et al., 2025b; Gonçalves et al., 2025). In this sense, the Grether-style exercise places the data in a language that is comparable to the traditional literature, while still recovering the same broad reduced-form distortions highlighted by the baseline model.
The individual-level estimates (see Figure 13) likewise show substantial heterogeneity, again in a way that is closely aligned with the baseline specification. Overinference is the most prevalent reduced-form pattern, followed by base-rate neglect, while underinference and updating against the signal are also present for a non-negligible subset of participants and base-rate overuse is essentially absent.
Figure 13: Number of times a bias is found to be significant ($p<0.05$) in the Grether-style model at the individual level
Appendix F Experiment instructions